### v2

@@ 1 Introduction
Previous work has aimed to address biases in AI models for medical imaging tasks by using generic bias mitigation methods. For example, Zong et al. (2022) proposed a framework with eleven algorithms which aimed to measure and mitigate biases in medical imaging classification datasets. However, the fairness gains achieved by such methods often cause reduced performance for some protected groups, a phenomenon known as the ‘fairness-accuracy trade-off’ Li and Li (2024). Furthermore, research has suggested that, whilst such generic bias mitigation approaches may appear to reduce bias when evaluated on internal validation sets, the fairness gains often do not hold when evaluated externally (Schrouff et al., 2022; Yang et al., 2024). We hypothesise that a possible reason for this lack of effectiveness is that these methods take a “blind” approach to addressing bias in AI models, i.e. they apply a generic bias mitigation algorithm that does not take account of the underlying cause of the bias. More recently, a new strand of research has emerged which attempts to understand the bias, with a view to using this knowledge to develop more informed mitigation approaches. For example, Glocker et al. (2023) analysed bias in chest X-ray classifiers, applying test-set resampling, multitask learning, and model inspection to provide insights into the way protected attributes were encoded in the AI model. Similarly, Olesen et al. (2025) used ‘slice discovery methods’ to reveal the root causes of sex bias, also in chest X-ray classifiers. Such methods potentially enable the development of bias mitigation techniques that are targeted at the underlying cause(s) of the bias, and could lead to improved and more robust mitigation, which will hold under the effects of domain shift to external validation sets.
@@ 4.1 Data
Previous work (Lee et al., 2022, 2023) has shown that bias is greater when the imbalance between ethnicities in the training set is greater. Therefore, we curated a dataset where biases would be significant to allow for better evaluation of mitigation methods. Our training set comprises 15 Black subjects and 4,221 White subjects, all randomly sampled from the full dataset. The remaining subjects were used as the internal validation test set. The demographic information of the subjects in the training and test sets can be seen in Table 1 and Table 2. Note that training was performed using only Black and White subjects but testing was performed using all other available subjects including Mixed, Asian, Chinese and Other. All ethnicity information refers to self-reported ethnicity based on the categories provided in the UK Biobank dataset.
@@
Table 1: Characteristics of subjects used in the training dataset. Mean (standard deviation) values are presented for each characteristic. Statistically significant differences between subject groups and the overall average are indicated with an asterisk * (p $<$ 0.05) and were determined using a two-tailed Student’s t-test.
@@
Table 2: Characteristics of subjects used in the internal validation test dataset. Mean (standard deviation) values are presented for each characteristic. One-way ANOVA was used to test for differences across ethnic groups. Statistically significant differences between groups are indicated with an asterisk. * (p $<$ 0.01).
@@
Table 3: Characteristics of subjects used in the external validation test set. Mean (standard deviation) values are presented for each characteristic. Statistically significant differences between subject groups and the overall average are indicated with an asterisk * (p $<$ 0.05) and were determined using a two-tailed Student’s t-test.
@@ 5.1 Internal validation
The internal validation results of the baseline and bias mitigation methods can be seen in Table 4 and Fig. 2. Oversampling was the only method to increase performance for the Black subjects such that there was no significant difference between the median DSC scores of the Black and White subjects. The method increased the median DSC for Black subjects by 0.045. Fairness performance metrics (SER and FG) also decreased, showing more equitable performance. Using oversampling also caused performance to increase for the other ethnicities compared to the baseline (see Fig. 2). This is perhaps surprising as performance on these other ethnicities was not optimised during training.
@@
Table 4: DSC and HD values for each of the bias mitigation methods tested on the internal validation set. The p-values were computed between White and Black subjects based on a two-sided Mann Whitney U test on the DSC scores. * p$<$ 0.05. The best median DSC score for Black and White subjects is shown in bold.
@@
Combining the mitigation methods did not produce significantly less biased results, as shown in Table 5. All three combinations decreased performance for White subjects but increased performance for Black subjects. The best combination was oversampling and Group DRO which reduced the performance for the White subjects the least and improved performance for the Black subjects the most, leading to the lowest FG.
Table 5: DSC and HD values for the combined bias mitigation methods tested on the internal validation set using original sized images. The p-values were computed between White and Black subjects based on a two-sided Mann Whitney U test on the DSC scores. * p$<$ 0.05. The best median DSC score for Black and White subjects is shown in bold.
@@
Table 7: DSC and HD values for each of the bias mitigation methods using the cascaded cropping approach on the internal validation data. The p-values were computed between White and Black subjects based on a two-sided Mann Whitney U test on the DSC scores. * p$<$ 0.05. The best median DSC score for Black and White subjects is shown in bold.
@@ 5.2 External validation
Table 8: DSC and HD values for the baseline and oversampling methods tested on the external validation set. The p-values were computed between White and Black subjects based on a two-sided Mann Whitney U test on the DSC scores. * p$<$ 0.05. The best median DSC score for Black and White subjects is shown in bold.
@@
The results from using the cascaded cropping based approach can be seen in Table 9. Tables S4 to S6 show the results broken down by cardiac region for the cascaded approach. Using this approach, performance was lower than the upper bound of using ground truth segmentations for cropping for both groups but performance remained higher than for the baseline model using uncropped images seen in Table 8. Performance increased for both groups when using oversampling and the FG decreased. Table S7 shows the results for the oversampling cascaded model broken down by sex and age.
Table 9: DSC and HD values for the models using cropping on the external validation dataset. The p-values were computed between White and Black subjects based on a two-sided Mann Whitney U test on the DSC scores. * p$<$ 0.05. The best median DSC score for Black and White subjects is shown in bold.
@@ 6.2 External validation
The best method overall (cascaded cropping used with oversampling) was based on an understanding of the root cause of bias in AI-based CMR analysis as reported in Lee et al. (2025a). This method resulted in not only reduced bias compared to the baseline model but also improved performance for both groups. This important result highlights that, when using a bias understanding-informed mitigation approach, there was no fairness-accuracy trade-off. The fairness-accuracy trade-off is sometimes used as a reason not to implement bias mitigation algorithms in clinical practice, due to the medical principle of non-maleficence. Our work suggests that this trade-off is not inevitable, but that it can be avoided by a careful analysis of bias in individual scenarios rather than application of a generic mitigation algorithm.
@@ 6.3 Limitations and future work
This work has some limitations. For example, only Black and White subjects were used during training. Future work could investigate the effect of mitigating biases using multiple ethnicities, in other protected attributes such as age and socioeconomic status, or in intersectional groups. There was also a relatively small number of subjects in the external dataset when compared to the number of subjects in the training and internal validation sets from the UK Biobank, so externally validating on a larger, more diverse dataset would be beneficial.
@@ 7 Conclusions
Table S1: Metrics for LVBP segmentation based on inference performed on external validation set using different models trained on UK Biobank data. The p-values were computed between White and Black subjects based on a two-sided Mann Whitney U test on the DSC scores. * p$<$ 0.05. The best median DSC score for Black and White subjects is shown in bold.
@@
Table S2: Metrics for LVM segmentation based on inference performed on the external validation set using different models trained on UK Biobank data. The p-values were computed between White and Black subjects based on a two-sided Mann Whitney U test on the DSC scores. * p$<$ 0.05. The best median DSC score for Black and White subjects is shown in bold.
@@
Table S3: Metrics for RVBP segmentation based on inference performed on the external validation set using different models trained on UK Biobank data. The p-values were computed between White and Black subjects based on a two-sided Mann Whitney U test on the DSC scores. * p$<$ 0.05. The best median DSC score for Black and White subjects is shown in bold.
@@
Table S4: Metrics for LVBP segmentation based on inference performed on the external validation set using the cascaded approach to crop the images. The p-values were computed between White and Black subjects based on a two-sided Mann Whitney U test on the DSC scores. * p$<$ 0.05. The best median DSC score for Black and White subjects is shown in bold.
@@
Table S5: Metrics for LVM segmentation based on inference performed on external validation set using the cascaded approach to crop the images. The p-values were computed between White and Black subjects based on a two-sided Mann Whitney U test on the DSC scores. * p$<$ 0.05. The best median DSC score for Black and White subjects is shown in bold.
@@
Table S6: Metrics for RVBP segmentation based on inference performed on external validation set using the cascaded approach to crop the images. The p-values were computed between White and Black subjects based on a two-sided Mann Whitney U test on the DSC scores. * p$<$ 0.05. The best median DSC score for Black and White subjects is shown in bold.
@@
Table S7: DSC and HD values by sex and age for the cascaded approach using oversampling on the external validation set. The p-values were computed based on a two-sided Mann Whitney U test on the DSC scores. * p$<$ 0.05. FG was calculated using $DSC_{male}-DSC_{female}$ and $DSC_{Age<50}-DSC_{Age>50}$
Cropped baseline cascaded model p = 0.12 Cropped oversampling cascaded model p = 0.68 Female Male Age $<$ 50 Age $>$ 50 n 28 56 52 32 Median DSC 0.913 0.927 0.922 0.922 IQR DSC 0.032 0.036 0.032 0.031 Median HD (mm) 4.069 4.143 3.869 4.208 IQR HD (mm) 1.927 1.402 1.558 1.738 SER 1.193 1.004 Fairness gap 0.0141 -0.0003

### v3

@@ 1 Introduction
Previous work has aimed to address biases in AI models for medical imaging tasks by using generic bias mitigation methods. For example, Zong et al. (2022) proposed a framework with eleven algorithms which aimed to measure and mitigate biases in medical imaging classification datasets. However, the fairness gains achieved by such methods often cause reduced performance for some protected groups, a phenomenon known as the ‘fairness-accuracy trade-off’ (Li and Li, 2024). Furthermore, research has suggested that, whilst such generic bias mitigation approaches may appear to reduce bias when evaluated on internal validation sets, the fairness gains often do not hold when evaluated externally (Schrouff et al., 2022; Yang et al., 2024). We hypothesise that a possible reason for this lack of effectiveness is that these methods take a “blind” approach to addressing bias in AI models, i.e. they apply a generic bias mitigation algorithm that does not take account of the underlying cause of the bias. More recently, a new strand of research has emerged which attempts to understand the bias, with a view to using this knowledge to develop more informed mitigation approaches. For example, Glocker et al. (2023) analysed bias in chest X-ray classifiers, applying test-set resampling, multitask learning, and model inspection to provide insights into the way protected attributes were encoded in the AI model. Similarly, Olesen et al. (2025) used ‘slice discovery methods’ to reveal the root causes of sex bias, also in chest X-ray classifiers. Such methods potentially enable the development of bias mitigation techniques that are targeted at the underlying cause(s) of the bias, and could lead to improved and more robust mitigation, which will hold under the effects of domain shift to external validation sets.
@@ 4.1 Data
Previous work (Lee et al., 2022, 2023) has shown that bias is greater when the imbalance between ethnicities in the training set is greater. Therefore, we curated a dataset where biases would be significant to allow for better evaluation of mitigation methods. Our main training set comprises 15 Black subjects and 4,221 White subjects, all randomly sampled from the full dataset. The remaining subjects were used as the internal validation test set. We also investigate the effect of using different proportions of Black and White subjects in the training set in the Supplementary Material. The demographic information of the subjects in the training and test sets can be seen in Table 1 and Table 2. Note that training was performed using only Black and White subjects but testing was performed using all other available subjects including Mixed, Asian, Chinese and Other. All ethnicity information refers to self-reported ethnicity based on the categories provided in the UK Biobank dataset.
@@
Table 1: Characteristics of subjects used in the training dataset. Mean (standard deviation) values are presented for each characteristic. Statistically significant differences between subject groups and the overall average are indicated with an asterisk * (p << 0.05) and were determined using a two-tailed Student’s t-test.
@@
Table 2: Characteristics of subjects used in the internal validation test dataset. Mean (standard deviation) values are presented for each characteristic. One-way ANOVA was used to test for differences across ethnic groups. Statistically significant differences between groups are indicated with an asterisk. * (p << 0.05).
@@
Table 3: Characteristics of subjects used in the external validation test set. Mean (standard deviation) values are presented for each characteristic. Statistically significant differences between subject groups and the overall average are indicated with an asterisk * (p << 0.05) and were determined using a two-tailed Student’s t-test.
@@ 4.8 Evaluation
4.9 Code availability
The code to reproduce this work is available at https://github.com/tiarnaleeKCL/nnUNet-bias-mitigation.
@@ 5.1 Internal validation
The internal validation results of the baseline and bias mitigation methods can be seen in Table 4 and Fig. 2. Statistical tests were performed using Mann-Whitney U tests on the overall DSC scores. Oversampling was the only method to increase performance for the Black subjects such that there was no significant difference between the median DSC scores of the Black and White subjects. The method increased the median DSC for Black subjects by 0.045. Fairness performance metrics (SER and FG) also decreased, showing more equitable performance. Using oversampling also caused performance to increase for the other ethnicities compared to the baseline (see Fig. 2). This is perhaps surprising as performance on these other ethnicities was not optimised during training. Further results on the effect of different levels of oversampling, both without and without cascaded cropping, can be seen in Figure S2. Plots of performance (median DSC) against FG for these two approaches can be seen in Figure S3. The effect of using different proportions of Black and White subjects in the training set are reported in Figure S1. Additional experiments using Asian and White subjects can be seen in Table S9 .
@@
Table 4: DSC and HD values for each of the bias mitigation methods tested on the internal validation set. The p-values were computed between White and Black subjects based on a two-sided Mann Whitney U test on the DSC scores. * p<< 0.05. The best median DSC score for Black and White subjects is shown in bold.
@@
Combining the mitigation methods did not produce significantly less biased results, as shown in Table 5. Mann-Whitney U tests were performed to test the significance of differences between the overall DSC scores. All three combinations decreased performance for White subjects but increased performance for Black subjects. The best combination was oversampling and Group DRO which reduced the performance for the White subjects the least and improved performance for the Black subjects the most, leading to the lowest FG.
Table 5: DSC and HD values for the combined bias mitigation methods tested on the internal validation set using original sized images. The p-values were computed between White and Black subjects based on a two-sided Mann Whitney U test on the DSC scores. * p<< 0.05. The best median DSC score for Black and White subjects is shown in bold.
@@
Table 7: DSC and HD values for each of the bias mitigation methods using the cascaded cropping approach on the internal validation data. The p-values were computed between White and Black subjects based on a two-sided Mann Whitney U test on the DSC scores. * p<< 0.05. The best median DSC score for Black and White subjects is shown in bold.
@@ 5.2 External validation
Table 8: DSC and HD values for the baseline and oversampling methods tested on the external validation set. The p-values were computed between White and Black subjects based on a two-sided Mann Whitney U test on the DSC scores. * p<< 0.05. The best median DSC score for Black and White subjects is shown in bold.
@@
The results from using the cascaded cropping based approach can be seen in Table 9. Mann-Whitney U tests were used to compare the overall DSC scores. Furthermore, the errors in the sizes of the bounding boxes using the cascaded approach compared to those calculated using the ground-truth segmentations can be seen in Table S4. Tables S5 to S7 show the results broken down by cardiac region for the cascaded approach. Using this approach, performance was lower than the upper bound of using ground truth segmentations for cropping for both groups but performance remained higher than for the baseline model using uncropped images seen in Table 8. Performance increased for both groups when using oversampling and the FG decreased. Table S8 shows the results for the oversampling cascaded model broken down by sex and age.
Table 9: DSC and HD values for the models using cropping on the external validation dataset. The p-values were computed between White and Black subjects based on a two-sided Mann Whitney U test on the DSC scores. * p<< 0.05. The best median DSC score for Black and White subjects is shown in bold.
@@ 6.2 External validation
The best method overall (cascaded cropping used with oversampling) was based on an understanding of the root cause of bias in AI-based CMR analysis as reported in Lee et al. (2025a). Specifically, in Lee et al. (2025a) it was reported that the main source of the distributional shift between ethnicities (and hence the bias in segmentation performance) was outside the heart. The cascaded cropping approach is a simple but elegant approach to removing the source of the bias in a robust way but maintaining (and even improving) segmentation performance. This important result highlights that, when using a bias understanding-informed mitigation approach in the context of CMR segmentation, there is no fairness-accuracy trade-off. The fairness-accuracy trade-off is sometimes used as a reason not to implement bias mitigation algorithms in clinical practice, due to the medical principle of non-maleficence. Our work suggests that this trade-off is not inevitable, but that it can be avoided by a careful analysis of bias in individual scenarios rather than application of a generic mitigation algorithm.
@@ 6.3 Limitations and future work
This work has some limitations. For example, only two ethnicities (Black/White and Asian/White) were used during training in our experiments. Future work could investigate the effect of mitigating biases using multiple ethnicities, in other protected attributes such as age and socioeconomic status, or in intersectional groups. There was also a relatively small number of subjects in the external dataset when compared to the number of subjects in the training and internal validation sets from the UK Biobank, so externally validating on a larger, more diverse dataset would be beneficial. In future work we also plan to investigate more advanced baseline bias mitigation methods than the ones used in this work. However, it should be noted that the literature on bias mitigation in segmentation (e.g. Puyol-Antón et al. (2021); Benčević et al. (2024); Siddiqui et al. (2024)) is limited compared to the more extensive literature in classification problems. (R3.1) Finally, the different loss functions used for Group DRO compared to reweighing and oversampling may have resulted in different regularisation for the models so future work could investigate the use of a single loss function.
@@ 7 Conclusions
Metrics for regions of the heart using the external validation set
Table S1: Metrics for LVBP segmentation based on inference performed on external validation set using different models trained on UK Biobank data. The p-values were computed between White and Black subjects based on a two-sided Mann Whitney U test on the DSC scores. * p<< 0.05. The best median DSC score for Black and White subjects is shown in bold.
@@
Table S2: Metrics for LVM segmentation based on inference performed on the external validation set using different models trained on UK Biobank data. The p-values were computed between White and Black subjects based on a two-sided Mann Whitney U test on the DSC scores. * p<< 0.05. The best median DSC score for Black and White subjects is shown in bold.
@@
Table S3: Metrics for RVBP segmentation based on inference performed on the external validation set using different models trained on UK Biobank data. The p-values were computed between White and Black subjects based on a two-sided Mann Whitney U test on the DSC scores. * p<< 0.05. The best median DSC score for Black and White subjects is shown in bold.
@@
Difference in the size of the bounding box between images
Table S4: Difference in the size of the bounding box between images cropped using the ground truth and cascaded approach for the internal set.
Experiment Error $x$ (%) Error $y$ (%)
Baseline 3.70 6.57
Oversampling 0.0 7.30
Reweighing 1.48 8.76
Group DRO 0.0 0.730
Metrics for regions of the heart for the external validation set using the cascaded cropping method
Table S5: Metrics for LVBP segmentation based on inference performed on the external validation set using the cascaded approach to crop the images. The p-values were computed between White and Black subjects based on a two-sided Mann Whitney U test on the DSC scores. * p<< 0.05. The best median DSC score for Black and White subjects is shown in bold.
@@
Table S6: Metrics for LVM segmentation based on inference performed on external validation set using the cascaded approach to crop the images. The p-values were computed between White and Black subjects based on a two-sided Mann Whitney U test on the DSC scores. * p<< 0.05. The best median DSC score for Black and White subjects is shown in bold.
@@
Table S7: Metrics for RVBP segmentation based on inference performed on external validation set using the cascaded approach to crop the images. The p-values were computed between White and Black subjects based on a two-sided Mann Whitney U test on the DSC scores. * p<< 0.05. The best median DSC score for Black and White subjects is shown in bold.
@@
DSC and HD values by sex and age for the cascaded approach
Table S8: DSC and HD values by sex and age for the cascaded approach using oversampling on the external validation set. The p-values were computed based on a two-sided Mann Whitney U test on the DSC scores. * p<< 0.05. FG was calculated using $DSC_{male}-DSC_{female}$ and $DSC_{Age<50}-DSC_{Age>50}$
Cropped baseline cascaded model p = 0.12 Cropped baseline cascaded model p = 0.68 Female Male Age << 50 Age >> 50 n 28 56 52 32 Median DSC 0.913 0.927 0.922 0.922 IQR DSC 0.032 0.036 0.032 0.031 Median HD (mm) 4.069 4.143 3.869 4.208 IQR HD (mm) 1.927 1.402 1.558 1.738 SER 1.193 1.004 Fairness gap 0.0141 -0.0003
Metrics for the bias mitigation methods using White and Asian subjects
Table S9: DSC and HD values for each of the bias mitigation methods tested on the internal validation set. The p-values were computed between White and Asian subjects based on a two-sided Mann Whitney U test on the DSC scores. * p<< 0.05. The best median DSC score for Asian and White subjects is shown in bold.
Baseline * Oversampling p = 0.17 Reweighing * Group DRO * Cropped baseline cascaded model Cropped oversampling cascaded model White Asian White Asian White Asian White Asian White Asian White Asian Median DSC 0.901 0.851 0.8990 0.9039 0.8971 0.8349 0.8971 0.8704 0.9584 0.9362 0.9576 0.8949 IQR DSC 0.0452 0.0635 0.0445 0.0678 0.0501 0.0686 0.0471 0.0584 0.0559 0.0502 0.0581 0.0583 Median HD (mm) 6.761 9.585 6.783 7.204 7.051 10.057 6.688 8.628 3.957 7.254 3.862 8.093 IQR HD (mm) 3.690 4.935 3.715 3.662 4.140 5.076 3.744 4.310 2.470 5.115 2.321 5.491 SER 1.498 1.051 1.604 1.260 1.533 2.477 Fairness gap 0.0494 -0.0049 0.0621 0.0267 0.0222 0.0627
The effect of changing the number of Black subjects in the training dataset
Figure S1: The effect of changing the number of Black subjects in the training dataset.
The effect of changing the level of oversampling in the training dataset
Figure S2: The effect of changing level of oversampling of Black subjects in the training dataset.
Median DSC against fairness gap for different levels of oversampling
Figure S3: Median DSC against fairness gap for different levels of oversampling in experiments using oversampling and cascaded cropping + oversampling. The fairness gap is calculate by subtracting the median DSC for Black subjects from the median DSC for White subjects.